There are two value of time for which s projectile  is at the same height .the sum of these two times is equal to
 

Grade:11

1 Answers

Eshan
5 years ago
Dear student,

Let the vertical compoenent of initial velocity beu.
HenceH=ut-\dfrac{1}{2}gt^2
\implies t^2-\dfrac{2u}{g}t+\dfrac{2H}{g}=0
Hence the sum of its roots is\dfrac{2u}{g}